Three capacitors of capacitances C1 = 10 nF, C2 = 15 nF, and C3 = 20 nF are connected in series to a 5-V battery. What is the charge stored in C2 and voltage across C3?


1
Expert's answer
2021-05-17T12:20:21-0400

"\frac{1}{C_T} = \frac{1}{C_1}+\frac{1}{C_2}+\frac{1}{C_3}"

"\frac{1}{C_T} = \frac{1}{10}+\frac{1}{15}+\frac {1}{20} =\frac{13}{60}"

"{C_T} = 4.6nF"

"Q_{C2}= V_{C2}\u00d7C_2"

"V_{C2}=V_S×\frac{C_T}{C_2}"

"V_{C2}=5×(\frac{4.6}{15})= 1.5V"


"V_{C3}=V_S×\frac{C_T}{C_3}"

"V_{C3}=5×(\frac{4.6}{20})= 1.15V"